# Lindbladian approach to $P(E)$ theory

### v1

@@ Abstract
The $P(E)$ theory describes inelastic tunneling of particles through junctions integrated into an electric circuit. The inelasticity stems from the interaction between the photonic field of the circuit and the tunneling particles possessing electric charge. In the conventional approach to the $P(E)$ theory, the tunneling rate and the electric current through the junction are derived using Fermi’s golden rule. The derivation is carried out to leading order in the tunnel coupling between the leads and includes an averaging over the environmental photonic degrees of freedom. Tracing out the environmental degrees of freedom is also commonly used for open quantum systems when deriving dynamical equations such as the Lindbladian. In this work, we reveal the relations between the $P(E)$ theory and the Lindbladian dynamics. We demonstrate that the assumptions of the Fermi’s golden rule are essentially equivalent with the Born-Markovian approximation used in the Lindbladian formalism. The resulting quantum master equation enables us to obtain not only the electric current but various other quantities, including for instance the heat current parametrized by the same $P(E)$ function, in a systematic and convenient way.
@@ I Introduction
Tunneling is one of the most peculiar quantum phenomena. In contrast to classical behavior, quantum transmission is also possible through potential barriers exceeding the total energy of the particle. Quantum tunneling has attracted much interest due to the wide range of applications including superconducting quantum interference devices [1, 2], scanning tunneling microscopy [3, 4, 5, 6, 7], atomic scale devices [8, 9, 10], quantum dots [11, 12, 13] and resonant tunneling diodes [14, 15].
@@
In most situations, the tunneling event can be considered a closed (i.e., unitary) process in the sense that the particle does not interact with the environment. In some experiments [16, 17, 6, 18, 19], however, the coupling to the electromagnetic environment significantly influences the transport properties. The electromagnetic field arises from the electric circuit into which the junction is embedded and typically plays an important role in junctions with low capacity (large charging energy). The effects of the coupling between a tunneling particle and the quantized electromagnetic (photonic) bath is conventionally described by the $P(E)$ theory [20, 21, 22, 23, 24]. The theory is based on Fermi’s golden rule describing the transmission probability rate between an initial and a final state with equal energies. By taking into account the energy of both the tunneling particles and the photons, and by averaging over the photonic degrees of freedom, one can derive an analytical expression for the $I(V)$ characteristics. The key quantity appearing in the formula is the $P(E)$ function which describes the probability that the tunneling particle emits a photon of energy $E$ to the environment.
Another very commonly used technique to describe the system-bath interaction is the Lindbladian formalism [25, 26, 27, 28, 29] which appears quite different in form and spirit: while the $P(E)$ theory describes inelastic tunneling of particles, the more general Lindbladian equation describes the dynamics of open quantum systems. The relation between the $P(E)$ theory and the Lindbladian dynamics has not been discussed so far. In this paper, we present the Lindbladian approach by applying the Born-Markovian approximation and by integrating out the photonic degrees of freedom. We also demonstrate how the resulting Lindblad equation can be used to calculate physical observables. We emphasize that the Lindbladian approach does not supersede the $P(E)$ approach; the aim of the present work is rather to create a link between the two theoretical fields. In particular, we compare the assumptions of the conventional Fermi’s golden rule and the Lindbladian approach. We show that the approximations of the two approaches are essentially the same. We believe, however, that the Lindbladian approach has one significant technical advantage: it provides a convenient framework for calculating various quantities beyond the $I(V)$ characteristics.
@@
The paper is structured as follows. After defining the problem in Sec. II, in Sec. III we review the main steps of the conventional approach by focusing on the approximations implied in the Fermi’s golden rule. These approximations are compared to the assumptions made during the derivation of the Lindbladian equation in Sec. IV. In Secs. V and VI, the electric and heat current are calculated using the Lindbladian framework.
@@ II Problem statement
$H=H_{B}+H_{0}+H_{T}.$ (1)
Here
@@
describes the electrons of the left ($L$) and the right ($R$) leads, $c_{Lk\sigma}$ and $c_{Rq\sigma}$ are the annihilation operators of electrons with wavenumber $k/q$ and spin $\sigma$. The energy spectra $\varepsilon_{L}(k)$ and $\varepsilon_{R}(q)$ will be kept general for most of the calculation.
Figure 1: Junction embedded into an electric circuit. Orange arrows indicate the possible tunneling processes through the insulating layer (grey). On the two sides of the junction, the left and the right leads are characterized by the energy spectra $\varepsilon_{L}(k)$ and $\varepsilon_{R}(q)$ and the chemical potentials of electrons, $\mu_{L/R}$. The electric circuit is characterized by the impedance of $Z(\omega)$, the bias voltage is $V$ and the junction itself has capacitance $C$.
@@
where $e^{-i\varphi}$ is a charge counting operator acting on the photonic field. From the circuit point of view, the junction is a capacitor. The operator $e^{-i\varphi}$ shifts the charge on the capacitor by the unit charge $e$. The tunneling amplitudes $T_{kq}$ include the factor of $1/\sqrt{N_{L}N_{R}}$ where $N_{L}$ and $N_{R}$ designate the number of momentum modes in the left and right lead, respectively.
@@ III Fermi’s golden rule approach
We briefly revisit the conventional derivation of the $P(E)$ theory for a single junction [20]. We will particularly focus on the approximations made in the derivation. These will later be compared to those of the Lindbladian approach.
@@
The $P(E)$ approach is based upon Fermi’s golden rule which is used to calculate the transmission rate. The transmission from the left lead state with energy $\varepsilon_{L}(k)$ to the right lead state with energy $\varepsilon_{R}(q)$ is characterized by the rate of
@@
$w_{k\sigma,B\rightarrow q\sigma,B^{\prime}}$ $=\frac{2\pi}{\hbar}\left|\langle k\sigma,B|T_{kq}c_{Rq\sigma}^{+}c_{Lk\sigma}e^{-i\varphi}|q\sigma,B^{\prime}\rangle\right|^{2}\times$ (4)
@@
×$\times\delta\left(\varepsilon_{k}+E_{B}-\varepsilon_{q}-E_{B^{\prime}}\right),$
where $B$ denotes the state of the photonic bath with energy $E_{B}$ and $\delta(\epsilon)$ is the Dirac delta. It is important to note that Fermi’s golden rule is obtained from the first-order perturbation theory and is valid if the tunneling is weak. To be more precise, the transition probability from the left lead state $|k\sigma,B\rangle$ to the right lead state $|q\sigma,B^{\prime}\rangle$ is obtained as $P_{k\sigma,B\rightarrow q\sigma,B^{\prime}}(t)=|\langle q\sigma,B^{\prime}|\Psi(t)\rangle|^{2}$ where $|\Psi(t)\rangle$ is the time evolved state with the initial condition of $|k\sigma,B\rangle$. In the matrix element $\langle q,B^{\prime}\sigma|\Psi(t)\rangle$, we keep only first order terms in the tunneling amplitude.
The derivation of Fermi’s golden rule includes an infinite-time limit of the average transmission probability, $w_{k\sigma,B\rightarrow q\sigma,B^{\prime}}=\lim_{t\to\infty}(P_{k\sigma,B\rightarrow q\sigma,B^{\prime}}(t)/t)$. This infinite time limit results in the decay of all transition probabilities between the initial and the final states with a finite energy difference. By using the infinite time limit, it is also assumed implicitly that the tunneling events are independent and that there is no temporal overlap between any two events.
@@
The overall transition rate from left to right is obtained as a sum of the products of the elementary transition probabilities $w_{k\sigma\rightarrow q\sigma}$ and occupation probabilities as described in Ref. [20]. It is assumed that the coupling between the photonic environment and the electron does not significantly affect the state of the environment. In reality, of course, the state of the environment is modified by the coupling but since the transition probabilities are already in second order in the tunneling these modifications result only in higher order corrections. For the same reason, the leads can also be assumed to remain in thermal equilibrium with the temperature $T$ and chemical potentials $\mu_{L}$ and $\mu_{R}$.
By taking advantage of the thermal equilibrium, the sum over the environmental degrees of freedom can be performed leading to the tunneling rate
@@
where we use the shorthand notation $f_{L}=f(\varepsilon_{L}(k)-\mu_{L})$ and $f_{R}=f(\varepsilon_{R}(q)-\mu_{R})$ where $f(\xi)=1/(e^{\beta\xi}+1)$ is the Fermi function and $\beta=1/(k_{\mathrm{B}}T)$ is the inverse temperature. The $P(E)$ function is defined as
$P(E)=\frac{1}{2\pi\hbar}\int_{-\infty}^{\infty}\mathrm{d}te^{\frac{i}{\hbar}Et}\mathrm{Tr}\left(\rho_{B}e^{i\varphi(t)}e^{-i\varphi(0)}\right)$ (6)
with $\rho_{B}$ the thermal equilibrium density operator of the photonic bath. The $P(E)$ function is related to the charge relaxation through the electric circuit and can be computed from the properties of the particular circuit [20]. From the electron point of view, the function $P(E)$ represents the probability that the tunneling electron emits a photon with an energy $E$ to the environment (for $E>0$) or absorb it (for $E<0$).
Similarly to Eq. (5), the total tunneling rate from right to left can also be obtained and the electric current is finally calculated as $I=e(\overrightarrow{\Lambda}-\overleftarrow{\Lambda})$.
We note that the derivation presented here differs from that of Ref. [20] in a canonical transformation and in that the bias voltage appears as the difference of the chemical potential in the two leads. The final result is the same. Our choice of presentation enables easier comparison between the Fermi golden rule and the Lindbladian approaches.
@@ IV Lindbladian approach
In this section, we present an alternative theoretical description of the coupling between the tunneling electrons and the electromagnetic field of the circuit. We consider the latter as an energy reservoir for the electrons and follow the standard procedure of the microscopic derivation of the Lindblad equation [27]. The goal of this section is to demonstrate that the approximations of the Fermi’s golden rule are essentially equivalent to the Born-Markovian approximation used for the Lindblad equation.
@@
In the standard procedure, the next assumption is that the environment relaxes much faster than the typical timescales of the system evolution. This Markovian approximation can also be rephrased as a statement that the time evolution of the system has no memory effects. Technically, the time argument $t^{\prime}$ of the density matrix $\rho_{S}(t^{\prime})$ is replaced by $t$ indicating that the system density matrix does not change essentially while the bath relaxes. In addition to this change of argument, we also set the lower limit of the integration in Eq. (7) to $\infty$- ∞ in accordance with the idea that no memory effects from the initial state are kept. The physical meaning of the Markovian assumption is that after a tunneling event (also referred to as a jump process in the Lindbladian language) the bath achieves complete relaxation before another event occurs. This is essentially the same assumption as the infinite-time limit applied in the Fermi’s golden rule approach.
@@
in the interaction picture. Here, we have defined the process operators $A_{p}(t)=A_{p}e^{-i\omega_{p}t}$ as
@@
where we have also introduced the composite process index $p=\{k,q,\sigma,j\}$. The operator $A_{p}$ describes a tunneling process between a left lead state with momentum $k$ and a right lead state with momentum $q$, spin $\sigma$ and $j$ is the direction of the tunneling. If the tunneling process occurs from left to right, $j=1$, and $j=-1$ for the opposite direction. In the exponent of Eq. (9), $j(p)$ indicates the direction of the process $p$. In the following, we will use $j$ and $j^{\prime}$ instead of $j(p)$ and $j(p^{\prime})$ for the sake of brevity.
@@
So far, we have derived the Lindblad equation for the tunneling particles which is the full dynamical equation for the density matrix. Based on the density matrix, one can compute various observables, not only the electric current but also, for example, the dissipated heat loss. We note that these quantities could possibly be computed using the conventional $P(E)$ approach as well but we believe that the Lindbladian approach provides a more convenient and systematic way.
Further advantage of the Lindbladian is that it provides a platform to add further effects to the model and also to describe slowly varying transient phenomena such us switching on the bias voltage. Furthermore, the wide range of numerical tools developed for solving Lindblad master equations can be applied [32].

### v2

# Inelastic tunneling through normal and superconducting junctions in the presence of photonic bath within Lindbladian formalism

@@ Abstract
An electron tunneling across a junction integrated into an electric circuit can generate an excitation in the photonic field (electromagnetic environment) and lose energy in the process. Such inelastic tunneling of particles is commonly described using the $P(E)$ theory. In the conventional approach to this theory, the tunneling rate and the electric current through the junction are derived using Fermi’s golden rule and by averaging over the environmental photonic degrees of freedom. In this work, we address the same problem of inelastic tunneling due to photonic environment in Lindbladian formalism and we present how the photonic degrees of freedom are traced out in quantum master equation approach. The resulting quantum master equation is parameterized by the same $P(E)$ function and enables us to obtain not only the electric current but various other quantities, for instance the heat current, in a systematic and convenient way. We also demonstrate that the Lindbladian formalism provides a comprehensive description of Bogoliubov quasiparticle tunneling through superconducting junctions and that it properly accounts for the coherence factors. The coherence factors become important if the normal-state density of states is particle-hole asymmetric.
@@ I Introduction
Tunneling is one of the most peculiar quantum phenomena. In contrast to classical behavior, quantum transmission is also possible through potential barriers exceeding the total energy of the particle. Quantum tunneling has attracted much interest due to the wide range of applications including superconducting quantum interference devices [1, 2], scanning tunneling microscopy [3, 4, 5, 6, 7], atomic scale devices [8, 9, 10], quantum dots [11, 12, 13], quantum optical devices [14] and resonant tunneling diodes [15, 16].
@@
In most situations, the tunneling event can be considered a closed (i.e., unitary) process in the sense that the particle does not interact with the environment. In some experiments [17, 18, 6, 19, 20], however, the coupling to the electromagnetic environment significantly influences the transport properties of charged particles. The electromagnetic field arises from the electric circuit into which the junction is embedded and typically plays an important role in junctions with low capacity (large charging energy). The effects of the coupling between a tunneling particle and the quantized electromagnetic (photonic) bath is conventionally described by the $P(E)$ theory [21, 22, 23, 24, 25]. The theory is based on Fermi’s golden rule describing the transmission probability rate between an initial and a final state with equal energies. By taking into account the energy of both the tunneling particles and the photons, and by averaging over the photonic degrees of freedom, one can derive an analytical expression for the $I(V)$ characteristics. The key quantity appearing in the formula is the $P(E)$ function which describes the probability that the tunneling particle emits a photon of energy $E$ to the environment.
Another very commonly used technique to describe the system-bath interaction is the Lindbladian formalism [26, 27, 28, 29, 30, 31] which appears quite different in form and spirit: while the $P(E)$ theory describes inelastic tunneling of particles, the more general Lindblad master equation describes the dynamics of open quantum systems. Tunneling phenomena have been extensively studied within the framework of quantum master equations [32, 33, 34, 35, 36]. The typical approach involves tracing out electronic degrees of freedom of the leads and the master equation is formulated for the degrees of freedom of an embedded system, such as a quantum dot.
In this paper, we present a different kind of Lindbladian approach. We integrate out only the photonic degrees of freedom, but keep all electronic degrees of freedom (including those in the leads) as system variables. We demonstrate how the resulting Lindblad equation is used to calculate physical observables. We emphasize that the Lindbladian approach does not supersede the $P(E)$ approach; the aim of the present work is rather to compare the conventional $P(E)$ theory to the Lindbladian approach and to demonstrate some benefits of the latter. In particular, we compare the assumptions behind the conventional Fermi’s golden rule and the Lindbladian approach. The relation between Fermi’s golden rule and classical Markovian master equations has been previously studied in the strict limit of infinitesimally small coupling in Ref. 37. In this work, we investigate what assumptions apply for small but finite coupling and show that these are essentially the same in both approaches. We believe, however, that the Lindbladian approach has one significant technical advantage: it provides an algorithmic procedure to obtain various quantities related to tunnel phenomena without the necessity of introducing further assumptions. This is demonstrated by considering the problem of Bogoliubov quasi-particle tunneling through superconducting junctions.
@@
The paper is structured as follows. After defining the problem in Sec. II, in Sec. III we review the main steps of the conventional approach by focusing on the approximations implied in the Fermi’s golden rule. These approximations are compared to the assumptions made during the derivation of the Lindbladian equation in Sec. IV. In Secs. V, the electric and heat current are calculated using the Lindbladian framework. In Sec. VI, we demonstrate how to apply the formalism to quasiparticle tunneling through superconducting junctions. In App. A we provide details on the calculation of the dissipated heat. In App. B we present the formalism for tracking the number of electrons in superconducting BCS states.
@@ II Problem statement
$H=H_{B}+H_{0}+H_{T},$ (1)
where $H_{B}$ and $H_{T}$ are the Hamiltonians describing the electromagnetic bath and the tunneling, while
@@
describes the electrons of the left ($L$) and the right ($R$) leads. Here $c_{Lk\sigma}$ and $c_{Rq\sigma}$ are the annihilation operators of electrons with wavenumber $k/q$ and spin $\sigma$. The dispersions $\varepsilon_{L}(k)$ and $\varepsilon_{R}(q)$ describe the energy spectra of the metallic leads and will be kept general for most of the calculation.
Figure 1: Junction embedded into an electric circuit. Orange arrows indicate the possible tunneling processes through the insulating layer (grey). On the two sides of the junction, the left and the right leads are characterized by the energy spectra $\varepsilon_{L}(k)$ and $\varepsilon_{R}(q)$ and the chemical potentials of electrons, $\mu_{L/R}$. The electric circuit is characterized by the impedance of $Z(\omega)$, the bias voltage is $V$ and the junction itself has a capacitance $C$.
@@
where $e^{-i\varphi}$ is a charge displacement operator acting on the photonic field. From the circuit point of view, the junction is a capacitor. The operator $e^{-i\varphi}$ shifts the charge on the capacitor by the unit charge $e$ [22, 21, 24]. The tunneling amplitudes $T_{kq}$ include the factor of $1/\sqrt{N_{L}N_{R}}$ where $N_{L}$ and $N_{R}$ designate the number of momentum modes in the left and right lead, respectively.
@@ III Fermi’s golden rule approach
We revisit the conventional derivation of the $P(E)$ theory for a single junction [21]. We will particularly focus on the approximations made in the derivation. These will later be compared to those of the Lindbladian approach.
@@
The $P(E)$ approach is based upon Fermi’s golden rule which is used to calculate the transmission rate. By using first order perturbation theory, the transmission from the left lead state with energy $\varepsilon_{L}(k)$ to the right lead state with energy $\varepsilon_{R}(q)$ is characterized by the probability of
@@
$p_{k\sigma,B\rightarrow q\sigma B^{\prime}}(t)=\frac{\left|\langle q\sigma,B^{\prime}|H_{T}|k\sigma,B\rangle\right|^{2}}{\hbar^{2}}\times$
@@
×$\times\left|\int_{0}^{t}e^{\frac{i}{\hbar}\left(\varepsilon_{L}(k)+E_{B}-\varepsilon_{R}(q)-E_{B^{\prime}}\right)t^{\prime}}\mathrm{d}t^{\prime}\right|^{2},$ (4)
where $B$ and $B^{\prime}$ denote the bath (photonic) degrees of freedom. In order to compute the total probability of the transition $k\sigma\rightarrow q\sigma$, one has to sum over the bath degrees of freedom:
$\begin{split}&p_{k\sigma\rightarrow q\sigma}(t)=\sum_{BB^{\prime}}\rho_{B}p_{k\sigma,B\rightarrow q\sigma,B^{\prime}}\\ &=\frac{|T_{kq}|^{2}}{\hbar^{2}}\int_{0}^{t}\mathrm{d}t_{1}\int_{0}^{t}\mathrm{d}t_{2}e^{\frac{i}{\hbar}(\varepsilon_{L}(k)-\varepsilon_{R}(q))(t_{1}-t_{2})}\tilde{P}(t_{1}-t_{2}),\end{split}$ (5)
where $\rho_{B}$ is the probability that the photonic bath is in the state $B$ and
$\tilde{P}(t)=\mathrm{Tr}_{B}\left[\hat{\rho}_{B}e^{i\varphi(t)}e^{-i\varphi(0)}\right]$ (6)
is the bath correlation function with $\hat{\rho}_{B}$ the density matrix of the electromagnetic environment. It is assumed that $\hat{\rho}_{B}$ describes a thermal equilibrium state with temperature $T$. In reality, the state of the environment is modified by the coupling but since the transition probabilities are already in second order in the tunneling these modifications result in higher order corrections only. The bath correlation function $\tilde{P}(t)$ can be characterized by several time scales, the longest of which will be denoted by $\tau_{B}$.
By substituting with the Fourier transform of the bath correlation function,
$P(E)=\frac{1}{2\pi\hbar}\int_{-\infty}^{\infty}\mathrm{d}te^{\frac{i}{\hbar}Et}\mathrm{Tr}\left(\hat{\rho}_{B}e^{i\varphi(t)}e^{-i\varphi(0)}\right),$ (7)
and integrating over the variables $t_{1}$ and $t_{2}$, the transmission probability is rewritten as
$\begin{split}&p_{k\sigma\rightarrow q\sigma}(t)=\\ &\frac{|T_{kq}|^{2}}{\hbar^{2}}\int_{-\infty}^{\infty}\mathrm{d}E\,P(E)F\left(\frac{\varepsilon_{L}(k)-\varepsilon_{R}(q)-E}{\hbar},t\right)\end{split}$ (8)
where
$F(\omega,t)=\frac{\sin^{2}\left(\frac{\omega t}{2}\right)}{\left(\frac{\omega}{2}\right)^{2}}.$ (9)
Note that $F(\omega,t)$ is a sharp function around $\omega=0$ if $t$ is large enough since it has its major contribution in the range of $\left(-\frac{2\pi}{t};\frac{2\pi}{t}\right)$. If the size of this interval is much smaller than the narrowest energy interval on which $P(E)$ changes significantly in the vicinity of $\varepsilon_{L}(k)-\varepsilon_{R}(q)$, the function $F(\omega,t)$ can be replaced by $2\pi t\delta(\omega)$, leading to what is effectively the Fermi’s golden rule. The narrowest such energy interval is determined by the longest characteristic time scale $\tau_{B}$, leading to the condition $\frac{2\pi}{t}\ll\frac{1}{\tau_{B}}$ or, dropping the factor of $2\pi$, to
$\tau_{B}\ll t.$ (10)
If the condition is fulfilled, the transmission probability is obtained as
$p_{k\sigma\rightarrow q\sigma}(t)=t\frac{2\pi|T_{kq}|^{2}}{\hbar}P\left(\varepsilon_{L}(k)-\varepsilon_{R}(q)\right)\,.$ (11)
The probability is valid if it does not reach too high values. By using the notation $R_{kq}=\frac{2\pi|T_{kq}|^{2}}{\hbar}P\left(\varepsilon_{L}(k)-\varepsilon_{R}(q)\right)$, the condition for the validity of the long time limit of the Fermi’s golden rule is given by
$\tau_{B}\ll t\ll R_{kq}^{-1}.$ (12)
Such a condition can be fulfilled by some range of $t$ if
$\tau_{B}\ll R_{kq}^{-1}$ (13)
holds. This means that the Fermi’s golden rule is valid if the bath relaxation time is much shorter than the inverse transition rate.
Note that this condition is in spirit the same as presented e.g. in Chapter 27.6 of Ref. [38], but modified due to the averaging over the bath degrees of freedom. If the conditions from Eq. (12) are fulfilled, the transmission rate for the $k\sigma\rightarrow q\sigma$ transition is finally obtained as
$w_{k\sigma\rightarrow q\sigma}=\partial_{t}p_{k\sigma\rightarrow q\sigma}(t)=\frac{2\pi|T_{kq}|^{2}}{\hbar}P\left(\varepsilon_{L}(k)-\varepsilon_{R}(q)\right).$ (14)
@@
The overall transition rate from left to right is obtained as a sum of the products of the elementary transition probabilities $w_{k\sigma\rightarrow q\sigma}$ and occupation probabilities. Since the transmission rate is already in leading order in the tunneling amplitude, the occupation probabilities in the leads can be assumed to equal their equilibrium value corresponding to the temperature $T$ and chemical potentials $\mu_{L}$ and $\mu_{R}$.
Hence, the overall tunneling rate is given by
@@
where we use the shorthand notation $f_{L}=f(\varepsilon_{L}(k)-\mu_{L})$ and $f_{R}=f(\varepsilon_{R}(q)-\mu_{R})$ with $f(\xi)=1/(e^{\beta\xi}+1)$ the Fermi function and $\beta=1/(k_{\mathrm{B}}T)$ the inverse temperature. The $P(E)$ function is related to the charge relaxation through the electric circuit and can be computed from the properties of the particular circuit [21]. From the electron point of view, the function $P(E)$ represents the probability that the tunneling electron emits a photon with energy $E$ to the environment (for $E>0$) or absorb it (for $E<0$).
Similarly to Eq. (15), the total tunneling rate from right to left can also be obtained and the electric current is finally calculated as $I=e(\overrightarrow{\Lambda}-\overleftarrow{\Lambda})$ as presented in Ref. 21.
@@ IV Lindbladian approach
In this section, we present an alternative theoretical description of the coupling between the tunneling electrons and the electromagnetic field of the circuit. We consider the latter as an energy reservoir for the electrons and follow the standard procedure of the microscopic derivation of the Lindblad equation [28]. One goal of this section is to compare the approximations of the Fermi’s golden rule to the Born-Markovian approximation used for the Lindblad equation. The derivation presented here involves tracing out photonic degrees of freedom only and keeping all electronic degrees of freedom as system variables in contrast to many previous theoretical works [32, 33, 34, 35, 36].
@@
In the standard procedure, the next assumption is that the environment relaxes much faster than the typical timescales of the system evolution. This Markovian approximation can also be rephrased as a statement that the time evolution of the system has no memory effects. Technically, the time argument $t^{\prime}$ of the density matrix $\rho_{S}(t^{\prime})$ is replaced by $t$ indicating that the system density matrix does not change essentially while the bath relaxes. In addition to the change of argument, we also set the lower limit of the integration in Eq. (16) to $\infty$- ∞ in accordance with the idea that no memory effects from the initial state are kept. These technical steps lead to the same result as temporal coarse-graining [39, 40] with the time interval which is much longer than the bath relaxation time $\tau_{B}$ and is much shorter than the typical timescale of changes in the system. The physical meaning of the Markovian assumption is that after a tunneling event (also referred to as a jump process in the Lindbladian language) the bath achieves complete relaxation before another event occurs. This is essentially the same assumption as the long time limit applied in the Fermi’s golden rule approach as formulated in Eqs. (12) and (13).
@@
in the interaction picture. Here, we have defined the process operators $A_{p}(t)=A_{p}e^{-i\omega_{p}t}$ as shown in Table 1.
@@
Table 1: Jump processes in the case of normal leads.
For the jump processes we have introduced the composite process index $p=\{k,q,\sigma,j\}$. The operator $A_{p}$ describes a tunneling process between a left lead state with momentum $k$ and a right lead state with momentum $q$, spin $\sigma$ and $j$ is the direction of the tunneling. If the tunneling process occurs from left to right, $j=1$, and $j=-1$ for the opposite direction. In the exponent of Eq. (18), $j(p)$ indicates the direction of the process $p$. In the following, we will use $j$ and $j^{\prime}$ instead of $j(p)$ and $j(p^{\prime})$ for the sake of brevity.
@@
So far, we have derived the Lindblad equation for the tunneling particles which is the full dynamical equation for the density matrix.